Grade 11Trigonometry

Particle is pojected from a horizontal plane(x-y) such that its velocitu vector is given by v=ai^ +(b-ct)j^ .ITs range from horizontal plane is given by?

kartikay kala
7 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
V=a i^+(b-ct)j^
Vx=Ux+axt
a=Ux+0
Ux=a 
x=Uxt+0
x=at----------------(1)

Vy=Uy+ayt
b-ct=Uy-gt
Uy=b and g=c
now,
y=Uyt+1/2ayt^2 
0 =bt-1/2ct^2
2b-ct=0
t=2b/c-----------(2)
put (2)in equation (1)
x=a2b/c=2ab/c